OpenStudy (anonymous):

A woman has 7 keys on a keyring, one of which fits the door she wants to unlock. She randomly selects a key tries it. If it does not unlock the door, she randomly selects another from these remaining and tries to unlock the door with it. She continues in this manner until the door is unlocked. Let X be the number of keys she tries before unlocking the door, counting the key that actually worked. Find the probability function of X. A woman has 7 keys on a keyring, one of which fits the door she wants to unlock. She randomly selects a key tries it. If it does not unlock the door, she randomly selects another from these remaining and tries to unlock the door with it. She continues in this manner until the door is unlocked. Let X be the number of keys she tries before unlocking the door, counting the key that actually worked. Find the probability function of X. @Mathematics

OpenStudy (anonymous):

isn't it just... pr(X = 1) = 1/7 pr(X = 2) = 1/6 pr(X = 3) = 1/5 pr(X = 4) = 1/4 pr(X = 5) = 1/3 pr(X = 6) = 1/2 pr(X = 7) = 1

OpenStudy (anonymous):

whoops... Pr(X = 1) = 1/7 Pr(X = 2) = 1/7*1/6 Pr(X = 3) = 1/7*1/6*1/5 Pr(X = 4) = 1/7*1/6*1/5*1/4 Pr(X = 5) = 1/7*1/6*1/5*1/4*1/3 Pr(X = 6) = 1/7*1/6*1/5*1/4*1/3*1/2 Pr(X = 7) = 1/7*1/6*1/5*1/4*1/3*1/2*1